The base of a 17-ft ladder is 7 feet away from the wall. How far above the floor is the top of the ladder? Give your answer to the nearest thousandth.
Answer by nerdybill(7384)

The base of a 17-ft ladder is 7 feet away from the wall. How far above the floor is the top of the ladder? Give your answer to the nearest thousandth.
.
Applying the Pythagorean theorem:
Let x = length above the floor
then
x^2 + 7^2 = 17^2
x^2 + 49 = 289
x^2 = 240
x = 15.492 feet